OpenStudy (k.binks):

Find the midpoint of PQ if P(-3, -4) and Q(5, -5)?

pooja195 (pooja195):

\[\huge~\rm~\frac{ x_2+x_1 }{ 2},\frac{ y_2+y_1 }{ 2 }\]

pooja195 (pooja195):

Mid point formula from here just insert numbers

OpenStudy (k.binks):

(-3+5)/2 and (-4-5)/2 = 2/2 and -9/2? So (1, -9/2)?

OpenStudy (mathmate):

yes! That's correct!

OpenStudy (k.binks):

thanks!